Grade 12Magnetism

A circular coil of one turn is formed by 6.28m length wire, which carries a current of 3.14A the magnetic field at the centre of coil is?Ans; 2×10^-6

Syed suha
8 Years agoGrade 12

3 Answers

Khimraj
8 Years ago
Radi
Khimraj
8 Years ago
Radius of coil is given by
2*3.14*r = 6.28
r = 1m.
Magnetic field at centre of cil is given by
B = uoI/2r
B = (4\pi*10-7*3.14)/(2*1)
B = 6.28\pi*10-7 Tesla.
Hope it clears If you like answer then please approve it.
Krishna Pawar
8 Years ago
I=3.14(π}